# RangeLDM: Fast Realistic LiDAR Point Cloud Generation

### v1

@@ 2 Related Work
LiDAR point clouds can be represented in various forms, encompassing raw point clouds, voxels, range views, and multi-view fusion. Point-based models [61, 84, 83, 63, 49, 75, 90] directly encode 3D objects from raw points using the PointNet [50] encoder and subsequently perform detection or segmentation based on point features. These methods wholly preserve the irregularity and locality of a point cloud but have relatively higher latency. Voxel-based methods [14, 35, 81, 93, 80, 34, 76, 62, 13, 85, 59, 66, 17] voxelize point clouds for convolutional neural networks to efficiently capture features. They are computationally effective but the desertion of fine-grained patterns degrades further refinement. Given that the range view is compact and compatible with the LiDAR sensor’s sampling process, range-image-based methods [18, 67, 8, 10, 36, 44, 70, 45, 79, 12, 33, 91, 2, 11] have been proposed to directly process range images for LiDAR perception. Multi-view fusion methods [60, 22, 26] amalgamate multiple representations, which yield better results at the expense of processing speed. In this paper, we adopt the range-view representation, which is congruent with the LiDAR sensor’s sampling process and can be efficiently encoded by image generation models.
Generative Models for Point Clouds.
@@
Given observed samples of interest, generative models aim to learn the underlying distribution of the data and generate new samples. The first 3D generative models [1, 64, 72] for point clouds are based on GANs [21]. They operate by generating point clouds and discriminating them from real samples in an adversarial manner. The second category of methods [40, 69, 47, 20, 19, 31, 46] employ VAEs [32] to explore the probabilistic latent space of 3D shapes. Auto-regressive models [4] have also been introduced to generate point clouds [68] from scratch or semantic contexts. PointFlow [82] and SoftFlow [30] leverage normalizing flows [52] to capture the likelihood of shapes and points. Notably, the recent success of denoising diffusion models (DDMs) [25] for image synthesis [54] has also been extended to the domain of 3D point cloud generation [41, 92, 7, 88, 48, 28]. However, these previous works often concentrate on the object level and are not well-suited for handling large scenes.
@@
In this paper, our focus lies in the generation of LiDAR point clouds. While [5] and [56] employ VAE or GANs for LiDAR point cloud generation, the realism achieved in their results is relatively limited. UltraLiDAR [78], on the other hand, utilizes VQ-VAE to generate voxelized LiDAR point clouds, but at the expense of introducing quantified losses and slow generation speed. LiDARGen [94] proposed a novel score-matching energy-based model to generate higher-quality LiDAR point clouds, but it samples very slowly and has degraded geometric details at a far range.
@@ 3 Background on Denoising Diffusion Models
$q\left(\mathbf{x}_{1:T}\mid\mathbf{x}_{0}\right):=$ $\prod_{t=1}^{T}q\left(\mathbf{x}_{t}\mid\mathbf{x}_{t-1}\right),$ (1)
$q\left(\mathbf{x}_{t}\mid\mathbf{x}_{t-1}\right):=$ $\mathcal{N}\left(\mathbf{x}_{t};\sqrt{1-\beta_{t}}\mathbf{x}_{t-1},\beta_{t}\mathbf{I}\right),$
where $\beta_{t}$ evolves over time according to a fixed or learnable schedule, which is structured such that the distribution of the final latent $q\left(\mathbf{x}_{T}\right)$ is a standard Gaussian. The joint distribution $p_{\theta}\left(\mathbf{x}_{0:T}\right)$, called the reverse diffusion process, is also defined as a Markov chain with Gaussian transition kernels using the mean $\mu_{\boldsymbol{\theta}}\left(\mathbf{x}_{t},t\right)$ and fixed variance $\rho_{t}$:
$p_{\boldsymbol{\theta}}\left(\mathbf{x}_{0:T}\right):=$ $p\left(\mathbf{x}_{T}\right)\prod_{t=1}^{T}p_{\boldsymbol{\theta}}\left(\mathbf{x}_{t-1}\mid\mathbf{x}_{t}\right),$ (2)
$p_{\boldsymbol{\theta}}\left(\mathbf{x}_{t-1}\mid\mathbf{x}_{t}\right):=$ $\mathcal{N}\left(\mathbf{x}_{t-1};\mu_{\boldsymbol{\theta}}\left(\mathbf{x}_{t},t\right),\rho_{t}^{2}\boldsymbol{I}\right).$
@@
$\min_{\theta}\mathbb{E}_{\mathbf{x},\boldsymbol{\epsilon}\sim\mathcal{N}(0,1),t}\left[\left\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta}\left(\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon},t\right)\right\|_{2}^{2}\right],\vspace{0pt}$ (3)
where we use the notation $\alpha_{t}:=1-\beta_{t}$ and $\bar{\alpha}_{t}:=\prod_{s=1}^{t}\alpha_{s}$.
@@ 4 The Proposed Method
Figure 2: The framework of the proposed RangeLDM. Firstly, we project point clouds onto high-quality range images via Hough Voting (Section 4.1). Subsequently, we train a VAE to compress the range images into low-dimensional latent features $\mathbf{z}_{0}$, which encodes the range images with the encoder $\mathcal{E}_{\zeta}$ and reconstructs range images from latent features with the decoder $\mathcal{G}_{\eta}$ (Section 4.2.1). Here, a range-guided discriminator $\mathcal{D}_{\tau}$ is introduced to guide the decoder in the reconstruction of 3D structures. We finally train a latent diffusion model to capture the distribution of the latent features (Section 4.2.2). With optional conditional inputs, the proposed method is applicable to tasks such as point cloud upsampling and inpainting (Section 4.3).
As shown in Fig. 2, we first project point clouds onto high-quality range images (as detailed in Section 4.1), considering the compact nature of the range view, which aligns seamlessly with the sampling process of LiDAR sensors. Our model training process is then focused on these projected 2D range images. We then compress range images into a low-dimensional latent space through a VAE (as detailed in Section 4.2.1). Subsequently, we proceed to train DDMs within the reduced-dimensional latent space (as discussed in Section 4.2.2). We discuss applications of the proposed model in Section 4.3, which involves unconditional generation and conditional generation. Implementation details are explained in Section 4.4.
@@ 4.1 High-Quality Range Projection
$r=\sqrt{x^{2}+y^{2}+z^{2}},\theta=\operatorname{atan}(y,x),\phi=\operatorname{atan}\left(z,\sqrt{x^{2}+y^{2}}\right).\vspace{-6pt}$ (4)
@@
$r=\sqrt{x^{2}+y^{2}+(z-h_{j})^{2}},\theta=\operatorname{atan}(y,x),\phi=\phi_{j},\vspace{-6pt}$ (5)
@@
We then rasterize points $(r,\theta,\phi)$ into a 2D cylindrical projection $R(u,v)$ (a.k.a., range image) of size $H\times W$ with
$u=\left(\left(\theta+\pi\right)/2\pi\right)W,v=j,\vspace{-6pt}$ (6)
where $(u,v)$ denotes the grid coordinate of a point in the range image. Thus, we obtain high-quality range images as illustrated at the bottom of Fig. 3. We denote the obtained range image as $\mathbf{x}\in\mathbb{R}^{H\times W\times 2}$, which comprises $H\times W$ pixels associated with both range and intensity $\{r,i\}$.
@@
Method $\text{MMD}_{\text{BEV}}$downarrow $\text{FRD}_{\text{range}}$downarrow $\text{JSD}_{\text{BEV}}$downarrow
@@ 4.2 Training
4.2.1 First-Stage: Dimensionality Reduction
In the first stage, we reduce the dimensionality of the obtained range images by a VAE, which removes imperceptible high-frequency details and leads to low-dimensional latent features that encode prominent information in the original range image.
@@ 4.2.1 First-Stage: Dimensionality Reduction
$\mathcal{L}_{ELBO}(\zeta,\eta)=\mathbb{E}_{q_{\zeta}(\mathbf{z}_{0}\mid\mathbf{x})}\left[\log p_{\boldsymbol{\eta}}(\mathbf{x}\mid\mathbf{z}_{0})\right]-\lambda_{1}D_{\mathrm{KL}}\left(q_{\boldsymbol{\zeta}}(\mathbf{z}_{0}\mid\mathbf{x})\|p(\mathbf{z}_{0})\right).\vspace{-6pt}$ (7)
@@
Range-Guided Discriminator
In order to mitigate the blurriness introduced by relying solely on the pixel-space $L_{1}$ reconstruction loss, we integrate the VAE with a patch-based adversarial discriminator [29, 15, 86, 54]. Furthermore, to ensure that the reconstruction remains confined within the manifold underlying the range image and exploit the geometric information from the spherical coordinates, we propose a range-guided discriminator $\mathcal{D}_{\tau}$, with parameters $\tau$. Specifically, we adapt the Meta-Kernel [18] to replace the standard convolution in the discriminator, aiming to learn convolution weights from relative spherical coordinates. The Meta-Kernel is formulated as
$\mathbf{h}_{i}^{\prime}=\mathcal{W}\left(\mathop{\mathcal{A}}\limits_{j\in\mathcal{N}(i)}\left(\Phi\left(\gamma(\mathbf{p}_{j},\mathbf{p}_{i})\right)\odot\mathbf{h}_{j}\right)\right),\vspace{-6pt}$ (8)
@@
$\begin{gathered}\gamma\left(\mathbf{p}_{j},\mathbf{p}_{i}\right):=\{r_{j}\cos(\Delta\theta)\cos(\Delta\phi)-r_{i},r_{j}\cos(\Delta\theta)\sin(\Delta\phi),r_{j}\sin(\Delta\theta)\},\\ \Delta\theta=\theta_{j}-\theta_{i},\quad\Delta\phi=\phi_{j}-\phi_{i}.\end{gathered}\vspace{-6pt}$ (9)
@@
$\zeta,\eta=\arg\mathop{\min}\limits_{\zeta,\eta}\mathop{\max}\limits_{\tau}\mathcal{L}_{ELBO}(\zeta,\eta)+\lambda_{2}\mathcal{L}_{adv}(\zeta,\eta,\tau),\vspace{-6pt}$ (10)
@@
4.2.2 Second-Stage: Latent Diffusion Modeling
In the second stage, we freeze the encoder and decoder of the learned VAE and train an LDM on the encoded latent feature $\mathbf{z}_{0}$ for distribution modeling of the range image. During the diffusion process, noise is added to the initial latent $\mathbf{z}_{0}$, resulting in a noisy latent $\mathbf{z}_{t}$ as in Eq. 1, where the noise level increases over time steps $t\in T$.
@@ 4.2.2 Second-Stage: Latent Diffusion Modeling
$L_{\text{LDM}}:=\mathbb{E}_{\mathcal{E}(x),\boldsymbol{\epsilon}\sim\mathcal{N}(0,1),t}\left[\left\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta}\left(z_{t},t\right)\right\|_{2}^{2}\right].\vspace{-6pt}$ (11)
@@
$L_{\text{LDM}}:=\mathbb{E}_{\mathcal{E}(x),\mathbf{y},\boldsymbol{\epsilon}\sim\mathcal{N}(0,1),t}\left[\left\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta}\left(\mathbf{z}_{t},t,\tau_{\theta}(\mathbf{y})\right)\right\|_{2}^{2}\right],\vspace{-6pt}$ (12)
@@ 4.3 Generation
Unconditional Generation.
With the decoder $\mathcal{G}_{\eta}$ and the LDM $\epsilon_{\theta}$, we establish a hierarchical generative model $p_{\eta,\theta}(\mathbf{x},\mathbf{z}_{0})=p_{\eta}(\mathbf{x}|\mathbf{z}_{0})p_{\theta}(\mathbf{z}_{0})$. That is, we first generate latent features with the LDM, and then map the latent features back to the original range image space using the decoder $\mathcal{D}_{\eta}$ for LiDAR point cloud generation.
Conditional Generation.
Given a condition $\mathbf{y}$, the conditional generative model is defined as $p_{\eta,\theta}(\mathbf{x},\mathbf{z}_{0}|\mathbf{y})=p_{\eta}(\mathbf{x}|\mathbf{z}_{0},\mathbf{y})p_{\theta}(\mathbf{z}_{0}|\mathbf{y})$. There could be various conditions for LiDAR point cloud generation, such as sparse point clouds, incomplete point clouds, camera or even text. Here, we illustrate two applications of conditional generation: LiDAR point cloud upsampling and inpainting.
(1)LiDAR Point Cloud Upsampling. This task takes a sparse point cloud [^1] Without causing ambiguity, the point clouds are converted to range images by default. $\mathbf{y}\in\mathbb{R}^{h\times W\times 2}$ as input (following the setting of LiDARGen) and expects the model to produce a denser counterpart $\hat{\mathbf{x}}\in\mathbb{R}^{H\times W\times 2}$, which is necessary for accurate LiDAR-based perception and density-insensitive domain adaptation. We train LDM following Eq. 12 with the sparse point cloud $\mathbf{y}$ and its ground-truth dense variant $\mathbf{x}$. The condition encoder $\tau_{\theta}$ is configured as a reshaper that transforms the sparse point cloud $\mathbf{y}$ into $\mathbf{y}^{\prime}\in\mathbb{R}^{h\times w\times 2f}$. The reshaped $\mathbf{y}^{\prime}$ is then concatenated with noisy latent variable $\mathbf{z}_{t}$ and fed into the UNet $\epsilon_{\theta}$ for noise prediction.

### v2

@@ 2 Related Work
LiDAR Representation. LiDAR point clouds can be represented in various forms, encompassing raw point clouds, voxels, range views, and multi-view fusion. Point-based models [62, 85, 84, 64, 49, 76, 91] directly encode 3D objects from raw points using the PointNet [50] encoder and subsequently perform detection or segmentation based on point features. These methods wholly preserve the irregularity and locality of a point cloud but have relatively higher latency. Voxel-based methods [14, 35, 82, 94, 81, 34, 77, 63, 13, 86, 60, 67, 17] voxelize point clouds for convolutional neural networks to efficiently capture features. They are computationally effective but the desertion of fine-grained patterns degrades further refinement. Given that the range view is compact and compatible with the LiDAR sensor’s sampling process, range-image-based methods [18, 68, 8, 10, 36, 44, 71, 45, 80, 12, 33, 92, 2, 11] have been proposed to directly process range images for LiDAR perception. Multi-view fusion methods [61, 22, 26] amalgamate multiple representations, which yield better results at the expense of processing speed. In this paper, we adopt the range-view representation, which is congruent with the LiDAR sensor’s sampling process and can be efficiently encoded by image generation models.
@@
Generative Models for Point Clouds. Given observed samples of interest, generative models aim to learn the underlying distribution of the data and generate new samples. The first 3D generative models [1, 65, 73] for point clouds are based on GANs [21]. They operate by generating point clouds and discriminating them from real samples in an adversarial manner. The second category of methods [40, 70, 47, 20, 19, 31, 46] employ VAEs [32] to explore the probabilistic latent space of 3D shapes. Auto-regressive models [4] have also been introduced to generate point clouds [69] from scratch or semantic contexts. PointFlow [83] and SoftFlow [30] leverage normalizing flows [53] to capture the likelihood of shapes and points. Notably, the recent success of denoising diffusion models (DDMs) [25] for image synthesis [55] has also been extended to the domain of 3D point cloud generation [41, 93, 7, 89, 48, 28]. However, these previous works often concentrate on the object level and are not well-suited for handling large scenes.
@@
In this paper, our focus lies in the generation of LiDAR point clouds. While [5] and [57] employ VAE or GANs for LiDAR point cloud generation, the realism achieved in their results is relatively limited. UltraLiDAR [79], on the other hand, utilizes VQ-VAE to generate voxelized LiDAR point clouds, but at the expense of introducing quantified losses and slow generation speed. LiDARGen [95] proposed a novel score-matching energy-based model to generate higher-quality LiDAR point clouds, but it samples very slowly and has degraded geometric details at a far range. A contemporary work LiDM [51] also utilized LDMs to generate LiDAR point clouds. However, while they focused on generating LiDAR point clouds conditional on multimodal data, we focus on improving the quality of point cloud generation directly.
@@ 3 Background on Denoising Diffusion Models
$\min_{\theta}\mathbb{E}_{\mathbf{x},\boldsymbol{\epsilon}\sim\mathcal{N}(0,1),t}\left[\left\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta}\left(\mathbf{x}_{t}+\boldsymbol{\epsilon},t\right)\right\|_{2}^{2}\right],$ (1)
with $t$ uniformly sampled from $\{1,...,T\}$.
@@ 4 The Proposed Method
Figure 2: The framework of the proposed RangeLDM. Firstly, we project point clouds onto high-quality range images via Hough Voting (Section 4.1). Subsequently, we train a VAE to compress the range images into low-dimensional latent features $\mathbf{z}_{0}$, which encodes the range images with the encoder $\mathcal{E}_{\zeta}$ and reconstructs range images from latent features with the decoder $\mathcal{G}_{\eta}$ (Section 4.2). Here, a range-guided discriminator $\mathcal{D}_{\tau}$ is introduced to guide the decoder in the reconstruction of 3D structures. We finally train a latent diffusion model to capture the distribution of the latent features (Section 4.2). With optional conditional inputs, the proposed method is applicable to tasks such as point cloud upsampling and inpainting (Section 4.3).
As shown in Fig. 2, we first project point clouds onto high-quality range images (as detailed in Section 4.1), considering the compact nature of the range view, which aligns seamlessly with the sampling process of LiDAR sensors. Our model training process is then focused on these projected 2D range images. We then compress range images into a low-dimensional latent space through a VAE (as detailed in Section 4.2). Subsequently, we proceed to train DDMs within the reduced-dimensional latent space (as discussed in Section 4.2). We discuss applications of the proposed model in Section 4.3, which involves unconditional generation and conditional generation. Implementation details are explained in Section 4.4.
@@ 4.1 High-Quality Range Projection
$r=\sqrt{x^{2}+y^{2}+z^{2}},\theta=\operatorname{atan}(y,x),\phi=\operatorname{atan}\left(z,\sqrt{x^{2}+y^{2}}\right).$ (2)
@@
$r=\sqrt{x^{2}+y^{2}+(z-h_{j})^{2}},\theta=\operatorname{atan}(y,x),\phi=\phi_{j},$ (3)
@@
We then rasterize points $(r,\theta,\phi)$ into a 2D cylindrical projection $R(u,v)$ (a.k.a., range image) of size $H\times W$ with $u=\left(\left(\theta+\pi\right)/2\pi\right)W,v=j,$ where $(u,v)$ denotes the grid coordinate of a point in the range image. Thus, we obtain high-quality range images as illustrated at the bottom of Fig. 3. We denote the obtained range image as $\mathbf{x}\in\mathbb{R}^{H\times W\times 2}$, which comprises $H\times W$ pixels associated with both range and intensity $\{r,i\}$.
@@
Method $\text{MMD}_{\text{BEV}}$downarrow FRD $\downarrow$ $\text{JSD}_{\text{BEV}}$downarrow
@@ 4.2 Training
First-Stage: Dimensionality Reduction. In the first stage, we reduce the dimensionality of the obtained range images by a VAE, which removes imperceptible high-frequency details and leads to low-dimensional latent features that encode prominent information in the original range image.
@@
$\mathcal{L}_{ELBO}(\zeta,\eta)=\mathbb{E}_{q_{\zeta}(\mathbf{z}_{0}\mid\mathbf{x})}\left[\log p_{\boldsymbol{\eta}}(\mathbf{x}\mid\mathbf{z}_{0})\right]-\lambda_{1}D_{\mathrm{KL}}\left(q_{\boldsymbol{\zeta}}(\mathbf{z}_{0}\mid\mathbf{x})\|p(\mathbf{z}_{0})\right).$ (4)
@@
Range-Guided Discriminator. In order to mitigate the blurriness introduced by relying solely on the pixel-space $L_{1}$ reconstruction loss, we integrate the VAE with a patch-based adversarial discriminator [29, 15, 87, 55]. Furthermore, to ensure that the reconstruction remains confined within the manifold underlying the range image and exploit the geometric information from the spherical coordinates, we propose a range-guided discriminator $\mathcal{D}_{\tau}$, with parameters $\tau$. Specifically, we adapt the Meta-Kernel [18] to replace the standard convolution in the discriminator, aiming to learn convolution weights from relative spherical coordinates. The Meta-Kernel is formulated as
$\mathbf{h}_{i}^{\prime}=\mathcal{W}\left(\mathop{\mathcal{A}}\limits_{j\in\mathcal{N}(i)}\left(\Phi\left(\gamma(\mathbf{p}_{j},\mathbf{p}_{i})\right)\odot\mathbf{h}_{j}\right)\right),$ (5)
@@
$\begin{gathered}\gamma\left(\mathbf{p}_{j},\mathbf{p}_{i}\right):=\{r_{j}\cos(\Delta\theta)\cos(\Delta\phi)-r_{i},r_{j}\cos(\Delta\theta)\sin(\Delta\phi),r_{j}\sin(\Delta\theta)\},\\ \Delta\theta=\theta_{j}-\theta_{i},\quad\Delta\phi=\phi_{j}-\phi_{i}.\end{gathered}$ (6)
@@
$\zeta,\eta=\arg\mathop{\min}\limits_{\zeta,\eta}\mathop{\max}\limits_{\tau}\mathcal{L}_{ELBO}(\zeta,\eta)+\lambda_{2}\mathcal{L}_{adv}(\zeta,\eta,\tau),$ (7)
@@
Second-Stage: Latent Diffusion Modeling In the second stage, we freeze the encoder and decoder of the learned VAE and train an LDM on the encoded latent feature $\mathbf{z}_{0}$ for distribution modeling of the range image. During the diffusion process, noise is added to the initial latent $\mathbf{z}_{0}$, resulting in a noisy latent $\mathbf{z}_{t}$, where the noise level increases over time steps $t\in T$.
@@
$L_{\text{LDM}}:=\mathbb{E}_{\mathcal{E}(x),\boldsymbol{\epsilon}\sim\mathcal{N}(0,1),t}\left[\left\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta}\left(z_{t},t\right)\right\|_{2}^{2}\right].$ (8)
@@
$L_{\text{LDM}}:=\mathbb{E}_{\mathcal{E}(x),\mathbf{y},\boldsymbol{\epsilon}\sim\mathcal{N}(0,1),t}\left[\left\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta}\left(\mathbf{z}_{t},t,\tau_{\theta}(\mathbf{y})\right)\right\|_{2}^{2}\right],$ (9)
@@ 4.3 Generation
Unconditional Generation. With the decoder $\mathcal{G}_{\eta}$ and the LDM $\epsilon_{\theta}$, we establish a hierarchical generative model $p_{\eta,\theta}(\mathbf{x},\mathbf{z}_{0})=p_{\eta}(\mathbf{x}|\mathbf{z}_{0})p_{\theta}(\mathbf{z}_{0})$. That is, we first generate latent features with the LDM, and then map the latent features back to the original range image space using the decoder $\mathcal{D}_{\eta}$ for LiDAR point cloud generation.
Conditional Generation. Given a condition $\mathbf{y}$, the conditional generative model is defined as $p_{\eta,\theta}(\mathbf{x},\mathbf{z}_{0}|\mathbf{y})=p_{\eta}(\mathbf{x}|\mathbf{z}_{0},\mathbf{y})p_{\theta}(\mathbf{z}_{0}|\mathbf{y})$. There could be various conditions for LiDAR point cloud generation, such as sparse point clouds, incomplete point clouds, camera or even text. Here, we illustrate two applications of conditional generation: LiDAR point cloud upsampling and inpainting.
LiDAR Point Cloud Upsampling. This task takes a sparse point cloud [^1] Without causing ambiguity, the point clouds are converted to range images by default. $\mathbf{y}\in\mathbb{R}^{h\times W\times 2}$ as input (following the setting of LiDARGen) and expects the model to produce a denser counterpart $\hat{\mathbf{x}}\in\mathbb{R}^{H\times W\times 2}$, which is necessary for accurate LiDAR-based perception and density-insensitive domain adaptation. We train LDM following Eq. 9 with the sparse point cloud $\mathbf{y}$ and its ground-truth dense variant $\mathbf{x}$. The condition encoder $\tau_{\theta}$ is configured as a reshaper that transforms the sparse point cloud $\mathbf{y}$ into $\mathbf{y}^{\prime}\in\mathbb{R}^{h\times w\times 2f}$. The reshaped $\mathbf{y}^{\prime}$ is then concatenated with noisy latent variable $\mathbf{z}_{t}$ and fed into the UNet $\epsilon_{\theta}$ for noise prediction.